# Resonances and resonance expansions for $\delta$-interactions on the half-space

### v2

@@ 5. Semiclassical Asymptotics of Resonances for the Half-Space
In general, as also the previous treatment shows, it is difficult to obtain explicit information on all resonances. In order to understand some properties of them, one can study their asymptotic behaviour on certain regimes. An example of these regimes is the high energy limit. To fix ideas we limit ourself to the one point interaction. This limit consist in studying the distribution of the poles of
$\Bigl{(}-\Delta_{\alpha,y}^{{\textup{D}}}-z^{2}\Bigr{)}^{-1},\quad\textup{for}\,\left\lvert z\right\rvert\to+\infty$
and the analogous distribution in the Neumann case $\Delta_{\alpha,y}^{{\textup{D}}}$. The high energy limit is an example of semiclassical limit. This is obtained by rescaling $z$ trough a reference constant $h>0$ (who mimics Planck’s constant) and passing to the limit $h\to 0$. In this way, the energy $\Bigl{(}\frac{z}{h}\Bigr{)}^{2}$ diverges, implying that the energies of the considered phenomena are well higher than the reference energy $h$. Correspondingly, we consider the behavior of the resolvent
@@
In applications, typically, the potential which describes the interaction (here the point interaction) is usually assumed to be energy dependent and hence dependent on $h$ in this setting. A common choice for the interaction is of the form $h^{-\beta}$ with $\beta>0$ (see for example [17] and [28]). We consider both positive and negative sign of the point charge. So the operator considered is $-h^{2}\Delta_{\pm h^{-\beta},y}^{\textup{D}}$ and its resonances are the poles in the complex lower half-plane of
@@
Here we study the half-space with both the signs of the delta interaction and include the Neumann b.c. at the end of the Section.
@@ 5.1. Dirichlet Boundary Condition
We start by studying the semiclassical resonance asymptotics for $-h^{2}\Delta_{\pm h^{-\beta},y}^{\textup{D}}$. The form of the resolvent implies that the equation for the resonances is
(5.1) ±$\pm h^{-\beta}-\frac{iz}{4\pi h}+\frac{e^{2iy_{3}\frac{z}{h}}}{8\pi y_{3}}=0,$
@@
$-e^{\pm 8\pi y_{3}h^{-\beta}}=\biggl{(}-2iy_{3}\frac{z}{h}\pm 8\pi y_{3}h^{-\beta}\biggr{)}e^{-2iy_{3}\frac{z}{h}\pm 8\pi y_{3}h^{-\beta}}.$
Setting $-w^{\pm}=-e^{\pm 8\pi y_{3}h^{-\beta}}$ and $x=-2iy_{3}\frac{z}{h}\pm 8\pi y_{3}h^{-\beta}$, the previous equation becomes $-w=xe^{x}$. The main difference in the two cases is that in the first $w\to+\infty$ as $h\to 0^{+}$, while in the second $w\to 0$.
The solutions $x$ in the complex plane of $-w=xe^{x}$ are a denumerable family denoted by $x=W_{k}(-w)$, where $W$ is the Lambert $W$ function (see [32]) and $k$ varies over $\mathbb{Z}$. Specifically we use that $W_{k}(w)$ can be expanded in a convergent series (note that this expansion is valid for both $w$ approaching zero and infinity)
@@
Here [${}{p}{q}$ ] are the Stirling numbers, which are the number of ways to arrange $p$ objects in $q$ cycles.
@@
The series Eq. 5.2 is absolutely convergent for $w$ large (or small) enough and $k\in\mathbb{Z}$. More precisely we have the tail estimate
(5.3) $\begin{split}\left\lvert R_{k}-\frac{\ln{\Bigl{(}\ln{(w)}+(2k+1)i\pi\Bigr{)}}}{\ln{(w)}+(2k+1)i\pi}\right\rvert\leq\sum_{j\geq 0,m\geq 1,(j,m)\neq(0,1)}\left\lvert c_{j,m}\frac{\ln^{m}{\Bigl{(}\ln{(w)}+(2k+1)i\pi\Bigr{)}}}{\Bigl{(}\ln{(w)}+(2k+1)i\pi\Bigr{)}^{j+m}}\right\rvert\leq 2\left\lvert\frac{\ln{\Bigl{(}\ln{(w)}+(2k+1)i\pi\Bigr{)}}}{\ln{(w)}+(2k+1)i\pi}\right\rvert^{2}\end{split}$
@@
(5.4) $\left\lvert\frac{\ln{\Bigl{(}\ln{(w)}+(2k+1)i\pi\Bigr{)}}}{\ln{(w)}+(2k+1)i\pi}\right\rvert\leq\frac{1}{2}.$
@@
We prove the Lemma for both $w$ big and small simultaneously. Inequality Eq. 5.4 is clearly fulfilled for $w$ sufficiently large (or small). To prove Eq. 5.3 we write
@@
(5.5) $S_{2}=\frac{1}{\left\lvert\ln{(w)}+(2k+1)i\pi\right\rvert}\sum_{m=1}^{+\infty}\left\lvert\frac{\ln{\Bigl{(}\ln{(w)}+(2k+1)i\pi\Bigr{)}}}{\ln{(w)}+(2k+1)i\pi}\right\rvert^{m}\sum_{j=1}^{+\infty}\frac{\left\lvert c_{j,m}\right\rvert}{\left\lvert\ln{(w)}+(2k+1)i\pi\right\rvert^{j-1}}.$
@@
Moreover, for large $w$, the first multiplicative term is decreasing in $w$ and each term in the sum is decreasing as well. Since all terms are positive, we can conclude that the expression above is decreasing for large $w$. This means that there is $N$ such that, setting $w=N$, $\forall m\geq 1$ we have
@@
In the same way we have that, for small $w$, the first multiplicative term is increasing in $w$ and each term in the sum is increasing and positive as well. So the whole expression is increasing for small $w$ and that means that ther is $N$ such that if $w=\frac{1}{N}$, $\forall m\geq 1$
@@
Given that $\left\lvert\ln{\Bigl{(}\ln{(w)}+(2k+1)i\pi\Bigr{)}}\right\rvert\to+\infty$ for both $w\to+\infty$ and $w\to 0$, we have that
@@
So we can write the resonances $z_{k}^{\pm}$ in terms of the Lambert $W$ function. For $-h^{2}\Delta_{h^{-\beta},y}^{\textup{D}}$ they are given by
(5.6) $\begin{split}z_{k}^{+}&=\frac{ih}{y_{3}}\biggl{(}W_{k}(-w)-8\pi y_{3}h^{-\beta}\biggr{)}\\ &=\frac{ih}{y_{3}}\biggl{(}\ln{(w)}+(2k+1)i\pi-\ln{\Bigl{(}\ln{(w)}+(2k+1)i\pi\Bigr{)}}-8\pi y_{3}h^{-\beta}+R_{k}\biggr{)}\\ &=\frac{ih}{y_{3}}\biggl{(}(2k+1)i\pi-\ln{\Bigl{(}\ln{(w)}+(2k+1)i\pi\Bigr{)}}+R_{k}\biggr{)},\end{split}$
where $k$ varies over $\mathbb{Z}$ and $w=e^{8\pi y_{3}h^{-\beta}}$. Instead, for $-h^{2}\Delta_{-h^{-\beta},y}^{\textup{D}}$ (here $w=e^{-8\pi y_{3}h^{-\beta}}$)
(5.7) $\begin{split}z_{k}^{-}&=\frac{ih}{y_{3}}\biggl{(}W_{k}(-w)+8\pi y_{3}h^{-\beta}\biggr{)}\\ &=\frac{ih}{y_{3}}\biggl{(}\ln{(w)}+(2k+1)i\pi-\ln{\Bigl{(}\ln{(w)}+(2k+1)i\pi\Bigr{)}}+8\pi y_{3}h^{-\beta}+R_{k}\biggr{)}\\ &=\frac{ih}{y_{3}}\biggl{(}(2k+1)i\pi-\ln{\Bigl{(}\ln{(w)}+(2k+1)i\pi\Bigr{)}}+R_{k}\biggr{)}\ .\end{split}$
@@
Let $\varepsilon\in(0,1)$ ve given. Then, for $k$ such that $z_{k}$ is given by Eq. 5.6 or Eq. 5.7 and
(5.8) $\varepsilon\leq\left\lvert z_{k}\right\rvert\leq 1/\varepsilon,$
@@
(5.9) $\frac{\varepsilon}{2}\leq\frac{\left\lvert k\right\rvert\pi h}{y_{3}}\leq\frac{2}{\varepsilon}.$
@@
Using Eq. 5.4 and the inverse triangular inequality we have that
@@
Combining this with Eq. 5.3 and Eq. 5.4 we get $\left\lvert\left\lvert R_{k}\right\rvert-\frac{1}{2}\right\rvert\leq\frac{1}{2}$, equivalent to $\left\lvert R_{k}\right\rvert\leq 1$. We now prove the first inequality in Eq. 5.9. For the sake of contradiction we suppose that $\frac{\left\lvert k\right\rvert\pi h}{y_{3}}<\frac{\varepsilon}{2}$. Combining the first inequality of Eq. 5.8, $\left\lvert R_{k}\right\rvert\leq 1$ and Eq. 5.6, we get
@@
To prove the second inequality Eq. 5.9 we suppose for the sake of contradiction that $\frac{\left\lvert k\right\rvert\pi h}{y_{3}}>\frac{2}{\varepsilon}$. By $\left\lvert R_{k}\right\rvert\leq 1$, Eq. 5.6 and triangular inequality we have that
@@
This, combined with the second inequality of Eq. 5.8, implies
@@
We now establish asymptotic behaviour for resonances apart from the ones near zero and near infinity. We distinguish between $10<\beta<$ 1 and $\beta>1$. This different treatment is needed because different values of $\beta$ changes the relative dominance between $\ln{(y)}$ and $\left\lvert 2k+1\right\rvert\pi$ in Eq. 5.6. We start with the case $\beta\in(0,1)$.
@@
Let $\beta\in(0,1)$ and $\varepsilon\in(0,1)$ be given. Then there is $h_{0}>0$ such that, when $h\in(0,h_{0}]$, all solutions of Eq. 5.1 satisfying $\varepsilon\leq\left\lvert z\right\rvert\leq{1}/{\varepsilon}$, obey
@@
Since $\beta\in(0,1)$, then the first inequality in Eq. 5.9 implies
(5.10) $\left\lvert(2k+1)i\pi\right\rvert\geq 2\left\lvert k\right\rvert\pi-\pi\geq\varepsilon y_{3}h^{-1}-\pi>\left\lvert\ln{(w)}\right\rvert=8\pi y_{3}h^{-\beta},$
@@
Using $\left\lvert R_{k}\right\rvert\leq 1$ and Eq. 5.10, we get
(5.11) $\left\lvert R^{\prime}_{k}\right\rvert\leq\left\lvert R_{k}\right\rvert+\left\lvert\ln{\left\lvert 1+\frac{\ln{(w)}}{(2k+1)i\pi}\right\rvert}\right\rvert+\left\lvert\operatorname{Arg}{\biggl{(}1+\frac{\ln{(w)}}{(2k+1)i\pi}\biggr{)}}\right\rvert\leq 1+\ln{2}+\frac{\pi}{2}\leq 4.$
@@
(5.12) $\begin{cases}\operatorname{Re}{z_{k}^{\pm}}=-\frac{(2k+1)\pi h}{2y_{3}}+\frac{h}{2y_{3}}\operatorname{Arg}{\Bigl{(}(2k+1)i\pi\Bigr{)}}-\frac{h}{2y_{3}}\operatorname{Im}{R^{\prime}_{k}}\\ \operatorname{Im}{z_{k}^{\pm}}=-\frac{h}{2y_{3}}\ln{\left\lvert(2k+1)\pi\right\rvert}+\frac{h}{2y_{3}}\operatorname{Re}{R^{\prime}_{k}}\end{cases}.$
We now separate real and imaginary part of Eq. 5.1
@@
$-4\frac{y_{3}}{h}\operatorname{Im}{z_{k}^{\pm}}=2\ln{\Bigl{(}2y_{3}h^{-1}\left\lvert\operatorname{Re}{z_{k}^{\pm}}\right\rvert\Bigr{)}}+\ln{(1+t)}\ \ \text{where}\ \ \ t=\frac{4y_{3}^{2}\Bigl{(}\operatorname{Im}{z_{k}^{\pm}}\Bigr{)}^{2}\pm 32\pi y_{3}^{2}h^{1-\beta}\operatorname{Im}{z_{k}^{\pm}}+64\pi^{2}y_{3}^{2}h^{2-2\beta}}{4y_{3}^{2}\Bigl{(}\operatorname{Re}{z_{k}^{\pm}}\Bigr{)}^{2}}.$
@@
(5.13) $0\leq-4\frac{y_{3}}{h}\operatorname{Im}{z_{k}^{\pm}}-2\ln{\Bigl{(}2y_{3}h^{-1}\left\lvert\operatorname{Re}{z_{k}^{\pm}}\right\rvert\Bigr{)}}\leq t.$
From Eq. 5.12 we have, using Eq. 5.9, Eq. 5.11 and $\operatorname{Arg}{\Bigl{(}(2k+1)i\pi\Bigr{)}}=\pi/2$, that, for sufficiently small $h$
@@
Plugging this inequality into Eq. 5.13 gives the desired inequality. ∎
A different behaviour happens for $\beta>1$.
@@
Let $\beta>1$ and $\varepsilon\in(0,1)$ be given. Then there is $h_{0}>0$ such that, when $h\in(0,h_{0}]$, all solutions to Eq. 5.1+ satisfying $\varepsilon\leq\left\lvert z\right\rvert\leq{1}/{\varepsilon}$ obey
@@
which, using Eq. 5.9, implies
(5.14) $\operatorname{Im}{z}+\frac{\varepsilon^{2}}{32y_{3}}h\ln{\Bigl{(}8\pi y_{3}h^{-\beta}\Bigr{)}}\Bigl{(}\operatorname{Re}{z}\Bigr{)}^{2}\leq\frac{1+96\varepsilon^{-4}}{4\pi y_{3}^{2}}h^{\beta+1}\ln{\Bigl{(}8\pi y_{3}h^{-\beta}\Bigr{)}}+\frac{\varepsilon^{-2}h^{2\beta-1}}{2\pi^{2}y_{3}}$
(5.15) $\operatorname{Im}{z}+\frac{8}{\varepsilon^{2}y_{3}}h\ln{\Bigl{(}8\pi y_{3}h^{-\beta}\Bigr{)}}\Bigl{(}\operatorname{Re}{z}\Bigr{)}^{2}\geq-\frac{1+96\varepsilon^{-4}}{4\pi y_{3}^{2}}h^{\beta+1}\ln{\Bigl{(}8\pi y_{3}h^{-\beta}\Bigr{)}}-\frac{\varepsilon^{-2}h^{2\beta-1}}{2\pi^{2}y_{3}}.$
@@
Since $\beta>1$, by the second inequality in Eq. 5.9, we have
@@
for $h$ small enough. Recalling Eq. 5.6, we write
@@
(5.16) $\begin{cases}\operatorname{Re}{z_{k}}=\frac{h}{2y_{3}}\Biggr{(}-(2k+1)\pi+\operatorname{Arg}{\biggr{(}1+\frac{i(2k+1)}{8y_{3}}h^{\beta}\biggr{)}}-\operatorname{Im}{R_{k}}\Biggr{)}\\ \operatorname{Im}{z_{k}}=\frac{h}{2y_{3}}\Biggr{(}-\ln{\Bigl{(}8\pi y_{3}h^{-\beta}\Bigr{)}}-\ln{\left\lvert 1+\frac{i(2k+1)}{8y_{3}}h^{\beta}\right\rvert}+\operatorname{Re}{R_{k}}\Biggr{)}\end{cases}.$
@@
with $t=\frac{(2k+1)^{2}}{64y_{3}^{2}}h^{2\beta}$. Considering the imaginary part in Eq. 5.16, we get
(5.17) $\left\lvert\operatorname{Im}{z_{k}}+\frac{h}{2y_{3}}\ln{\Bigl{(}8\pi y_{3}h^{-\beta}\Bigr{)}}\right\rvert=\frac{h}{2y_{3}}\left\lvert\operatorname{Re}{R_{k}}-\ln{\left\lvert 1+\frac{i(2k+1)}{8y_{3}}h^{\beta}\right\rvert}\right\rvert.$
@@
$\begin{split}\left\lvert\frac{\ln{\Bigl{(}\ln{(w)}+(2k+1)i\pi\Bigr{)}}}{\ln{(w)}+(2k+1)i\pi}\right\rvert\leq\frac{\left\lvert\ln{\left\lvert\ln{(w)}+(2k+1)i\pi\right\rvert}+i\frac{\pi}{2}\right\rvert}{\left\lvert 8\pi y_{3}h^{-\beta}+(2k+1)i\pi\right\rvert}=\frac{\ln{\Bigl{(}h^{-\beta}\Bigr{)}}}{h^{-\beta}}\frac{\left\lvert\ln{\left\lvert\ln{(w)}+(2k+1)i\pi\right\rvert}+i\frac{\pi}{2}\right\rvert}{\ln{\Bigl{(}h^{-\beta}\Bigr{)}}\left\lvert 8\pi y_{3}+(2k+1)i\pi h^{\beta}\right\rvert},\end{split}$

### v3

@@ 5. Semiclassical Asymptotics of Resonances for the Half-Space
In general, as also the previous treatment shows, it is difficult to obtain explicit information on all resonances. In order to understand some properties of them, one can study their asymptotic behavior on certain regimes. An example of these regimes is the high energy limit. To fix ideas we limit ourself to the single point interaction in the half-space. This limit consist in studying the distribution of the poles of
$\Bigl(-\Delta_{\alpha,y}^{{\textup{D}}}-z^{2}\Bigr)^{-1},\quad\textup{for}\,\left\lvert z\right\rvert\to$$\infty$
and the analogous distribution in the Neumann case $\Delta_{\alpha,y}^{{\textup{N}}}$. The high energy limit is an example of semiclassical limit. This is obtained by rescaling $z$ trough a reference constant $h>0$ (who mimics Planck’s constant) and passing to the limit $h\to 0$. In this way, the energy $\Bigl(\frac{z}{h}\Bigr)^{2}$ diverges, implying that the energies of the considered phenomena are well higher than the reference energy $h$. Correspondingly, we consider the behavior of the resolvent
@@
In applications, typically, the potential which describes the interaction (here the point interaction) is usually assumed to be energy dependent and hence dependent on $h$ in this setting. A common choice for the interaction is of the form $h^{-\beta}$ with $\beta>0$ (see for example [23] and [35]). We consider both positive and negative sign of the point charge. So the operator considered is $h^{2}\Delta_{\pm h^{-\beta},y}^{\textup{D}}$ and its resonances are the poles in the complex lower half-plane of
@@
Here we study the half-space with both the signs of the delta interaction and we include the results for the Neumann b.c. at the end of the Section.
@@ 5.1. Dirichlet Boundary Condition
We start by studying the semiclassical resonance asymptotics for $h^{2}\Delta_{\pm h^{-\beta},y}^{\textup{D}}$. The form of the resolvent implies that the equation for the resonances is
$\pm h^{-\beta}-\frac{iz}{4\pi h}+\frac{e^{2iy_{3}\frac{z}{h}}}{8\pi y_{3}}=0,$ (5.1)
@@
$e^{\pm 8\pi y_{3}h^{-\beta}}=\biggl(-2iy_{3}\frac{z}{h}\pm 8\pi y_{3}h^{-\beta}\biggr)e^{-2iy_{3}\frac{z}{h}\pm 8\pi y_{3}h^{-\beta}}.$
Setting $w^{\pm}=-e^{\pm 8\pi y_{3}h^{-\beta}}$ and $x=-2iy_{3}\frac{z}{h}\pm 8\pi y_{3}h^{-\beta}$, the previous equation becomes $w=xe^{x}$. The main difference in the two cases is that in the first $w\to+\infty$ as $h\to 0^{+}$, while in the second $w\to 0$.
The solutions $x$ in the complex plane of $w=xe^{x}$ are a denumerable family denoted by $x=W_{k}(-w)$, where $W$ is the Lambert $W$ function (see [21]) and $k$ varies over $\mathbb{Z}$. Specifically we use the fact that $W_{k}$ can be expanded in a convergent series (note that this expansion is valid for both $w$ approaching zero and infinity)
@@
[…]
@@
The series Eq.˜5.2 is absolutely convergent for $w$ large (or small) enough and $k\in\mathbb{Z}$. More precisely we have the tail estimate
$\begin{split}\left\lvert R_{k}-\frac{\ln{\Bigl(\ln{(w)}+(2k+1)i\pi\Bigr)}}{\ln{(w)}+(2k+1)i\pi}\right\rvert\leq\sum_{j\geq 0,m\geq 1,(j,m)\neq(0,1)}\left\lvert c_{j,m}\frac{\ln^{m}{\Bigl(\ln{(w)}+(2k+1)i\pi\Bigr)}}{\Bigl(\ln{(w)}+(2k+1)i\pi\Bigr)^{j+m}}\right\rvert\leq 2\left\lvert\frac{\ln{\Bigl(\ln{(w)}+(2k+1)i\pi\Bigr)}}{\ln{(w)}+(2k+1)i\pi}\right\rvert^{2}\end{split}$ (5.3)
@@
$\left\lvert\frac{\ln{\Bigl(\ln{(w)}+(2k+1)i\pi\Bigr)}}{\ln{(w)}+(2k+1)i\pi}\right\rvert\leq\frac{1}{2}.$ (5.4)
@@
We prove the Lemma for both $w$ large and small simultaneously. Inequality Eq.˜5.4 is clearly fulfilled for $w$ sufficiently large (or small). To prove Eq.˜5.3 we write
@@
$S_{2}=\frac{1}{\left\lvert\ln{(w)}+(2k+1)i\pi\right\rvert}\sum_{m=1}^{+\infty}\left\lvert\frac{\ln{\Bigl(\ln{(w)}+(2k+1)i\pi\Bigr)}}{\ln{(w)}+(2k+1)i\pi}\right\rvert^{m}\sum_{j=1}^{+\infty}\frac{\left\lvert c_{j,m}\right\rvert}{\left\lvert\ln{(w)}+(2k+1)i\pi\right\rvert^{j-1}}.$ (5.5)
@@
Moreover, for large $w$, the first multiplicative term is decreasing in $w$ and each term in the sum is decreasing as well. Since all terms are positive, we can conclude that the expression above is decreasing for large $w$. This means that there is $N$ such that, setting $w=N$, for all $m\geq 1$ we have
@@
In the same way we have that, for small $w$, the first multiplicative term is increasing in $w$ and each term in the sum is increasing and positive as well. So the whole expression is increasing for small $w$ and that means that ther is $N$ such that if $w=\frac{1}{N}$, for all $m\geq 1$
@@
Given that $\left\lvert\ln{\Bigl(\ln{(w)}+(2k+1)i\pi\Bigr)}\right\rvert\to$$\infty$ for both $w\to+\infty$ and $w\to 0$, we have that
@@
So we can write the resonances $z_{k}^{\pm}$ in terms of the Lambert $W$ function. For $h^{2}\Delta_{h^{-\beta},y}^{\textup{D}}$ they are given by
$\begin{split}z_{k}^{+}&=\frac{ih}{y_{3}}\biggl(W_{k}(-w)-8\pi y_{3}h^{-\beta}\biggr)\\ &=\frac{ih}{y_{3}}\biggl(\ln{(w)}+(2k+1)i\pi-\ln{\Bigl(\ln{(w)}+(2k+1)i\pi\Bigr)}-8\pi y_{3}h^{-\beta}+R_{k}\biggr)\\ &=\frac{ih}{y_{3}}\biggl((2k+1)i\pi-\ln{\Bigl(\ln{(w)}+(2k+1)i\pi\Bigr)}+R_{k}\biggr),\end{split}$ (5.6)
where $k$ varies over $\mathbb{Z}$ and $w=e^{8\pi y_{3}h^{-\beta}}$. Instead, for $h^{2}\Delta_{-h^{-\beta},y}^{\textup{D}}$ (here $w=e^{-8\pi y_{3}h^{-\beta}}$)
$\begin{split}z_{k}^{-}&=\frac{ih}{y_{3}}\biggl(W_{k}(-w)+8\pi y_{3}h^{-\beta}\biggr)\\ &=\frac{ih}{y_{3}}\biggl(\ln{(w)}+(2k+1)i\pi-\ln{\Bigl(\ln{(w)}+(2k+1)i\pi\Bigr)}+8\pi y_{3}h^{-\beta}+R_{k}\biggr)\\ &=\frac{ih}{y_{3}}\biggl((2k+1)i\pi-\ln{\Bigl(\ln{(w)}+(2k+1)i\pi\Bigr)}+R_{k}\biggr)\ .\end{split}$ (5.7)
@@
Let $\varepsilon\in(0,1)$ be given. Then, for $k$ such that $z_{k}$ is given by Eq.˜5.6 or Eq.˜5.7 and
$\varepsilon\leq\left\lvert z_{k}\right\rvert\leq 1/\varepsilon,$ (5.8)
@@
$\frac{\varepsilon}{2}\leq\frac{\left\lvert k\right\rvert\pi h}{y_{3}}\leq\frac{2}{\varepsilon}.$ (5.9)
@@
Using Eq.˜5.4 and the inverse triangular inequality we have that
@@
Combining this with Eq.˜5.3 and Eq.˜5.4 we get $\left\lvert\left\lvert R_{k}\right\rvert-\frac{1}{2}\right\rvert\leq\frac{1}{2}$, equivalent to $\left\lvert R_{k}\right\rvert\leq 1$. We now prove the first inequality in Eq.˜5.9. For the sake of contradiction we suppose that $\frac{\left\lvert k\right\rvert\pi h}{y_{3}}<\frac{\varepsilon}{2}$. Combining the first inequality of Eq.˜5.8, $\left\lvert R_{k}\right\rvert\leq 1$ and Eq.˜5.6, we get
@@
To prove the second inequality Eq.˜5.9 we suppose for the sake of contradiction that $\frac{\left\lvert k\right\rvert\pi h}{y_{3}}>\frac{2}{\varepsilon}$. By $\left\lvert R_{k}\right\rvert\leq 1$, Eq.˜5.6 and triangular inequality we have that
@@
This, combined with the second inequality of Eq.˜5.8, implies
@@
We now establish asymptotic behaviour for resonances apart from the ones near zero and near infinity. We distinguish between $0<\beta<1$ and $\beta>1$. This different treatment is needed because different values of $\beta$ changes the relative dominance between $\ln{(y)}$ and $\left\lvert 2k+1\right\rvert\pi$ in Eq.˜5.6. We start with the case $\beta\in(0,1)$.
@@
Let $\beta\in(0,1)$ and $\varepsilon\in(0,1)$ be given. Then there is $h_{0}>0$ such that, when $h\in(0,h_{0}]$, all solutions of Eq.˜5.1 satisfying $\varepsilon\leq\left\lvert z\right\rvert\leq{1}/{\varepsilon}$, obey
@@
Since $\beta\in(0,1)$, then the first inequality in Eq.˜5.9 implies
$\left\lvert(2k+1)i\pi\right\rvert\geq 2\left\lvert k\right\rvert\pi-\pi\geq\varepsilon y_{3}h^{-1}-\pi>\left\lvert\ln{(w)}\right\rvert=8\pi y_{3}h^{-\beta},$ (5.10)
@@
Using $\left\lvert R_{k}\right\rvert\leq 1$ and Eq.˜5.10, we get
$\left\lvert R^{\prime}_{k}\right\rvert\leq\left\lvert R_{k}\right\rvert+\left\lvert\ln{\left\lvert 1+\frac{\ln{(w)}}{(2k+1)i\pi}\right\rvert}\right\rvert+\left\lvert\operatorname{Arg}{\biggl(1+\frac{\ln{(w)}}{(2k+1)i\pi}\biggr)}\right\rvert\leq 1+\ln{2}+\frac{\pi}{2}\leq 4.$ (5.11)
@@
$\begin{cases}\operatorname{Re}{z_{k}^{\pm}}=-\frac{(2k+1)\pi h}{2y_{3}}+\frac{h}{2y_{3}}\operatorname{Arg}{\Bigl((2k+1)i\pi\Bigr)}-\frac{h}{2y_{3}}\operatorname{Im}{R^{\prime}_{k}}\\ \operatorname{Im}{z_{k}^{\pm}}=-\frac{h}{2y_{3}}\ln{\left\lvert(2k+1)\pi\right\rvert}+\frac{h}{2y_{3}}\operatorname{Re}{R^{\prime}_{k}}\end{cases}.$ (5.12)
We now separate real and imaginary part of Eq.˜5.1
@@
$4\frac{y_{3}}{h}\operatorname{Im}{z_{k}^{\pm}}=2\ln{\Bigl(2y_{3}h^{-1}\left\lvert\operatorname{Re}{z_{k}^{\pm}}\right\rvert\Bigr)}+\ln{(1+t)}\ \ \text{where}\ \ \ t=\frac{4y_{3}^{2}\Bigl(\operatorname{Im}{z_{k}^{\pm}}\Bigr)^{2}\pm 32\pi y_{3}^{2}h^{1-\beta}\operatorname{Im}{z_{k}^{\pm}}+64\pi^{2}y_{3}^{2}h^{2-2\beta}}{4y_{3}^{2}\Bigl(\operatorname{Re}{z_{k}^{\pm}}\Bigr)^{2}}.$
@@
$0\leq-4\frac{y_{3}}{h}\operatorname{Im}{z_{k}^{\pm}}-2\ln{\Bigl(2y_{3}h^{-1}\left\lvert\operatorname{Re}{z_{k}^{\pm}}\right\rvert\Bigr)}\leq t.$ (5.13)
From Eq.˜5.12 we have, using Eq.˜5.9, Eq.˜5.11 and $\operatorname{Arg}{\Bigl((2k+1)i\pi\Bigr)}=\pi/2$, that, for sufficiently small $h$
@@
Plugging this inequality into Eq.˜5.13 gives the desired inequality. ∎
A different behavior happens for $\beta>1$.
@@
Let $\beta>1$ and $\varepsilon\in(0,1)$ be given. Then there is $h_{0}>0$ such that, when $h\in(0,h_{0}]$, all solutions to Eq.˜5.1+ satisfying $\varepsilon\leq\left\lvert z\right\rvert\leq{1}/{\varepsilon}$ obey
@@
which, using Eq.˜5.9, implies
$\operatorname{Im}{z}+\frac{\varepsilon^{2}}{32y_{3}}h\ln{\Bigl(8\pi y_{3}h^{-\beta}\Bigr)}\Bigl(\operatorname{Re}{z}\Bigr)^{2}\leq\frac{1+96\varepsilon^{-4}}{4\pi y_{3}^{2}}h^{\beta+1}\ln{\Bigl(8\pi y_{3}h^{-\beta}\Bigr)}+\frac{\varepsilon^{-2}h^{2\beta-1}}{2\pi^{2}y_{3}}$ (5.14)
$\operatorname{Im}{z}+\frac{8}{\varepsilon^{2}y_{3}}h\ln{\Bigl(8\pi y_{3}h^{-\beta}\Bigr)}\Bigl(\operatorname{Re}{z}\Bigr)^{2}\geq-\frac{1+96\varepsilon^{-4}}{4\pi y_{3}^{2}}h^{\beta+1}\ln{\Bigl(8\pi y_{3}h^{-\beta}\Bigr)}-\frac{\varepsilon^{-2}h^{2\beta-1}}{2\pi^{2}y_{3}}.$ (5.15)
@@
Since $\beta>1$, by the second inequality in Eq.˜5.9, we have
@@
for $h$ small enough. Recalling Eq.˜5.6, we write
@@
$\begin{cases}\operatorname{Re}{z_{k}}=\frac{h}{2y_{3}}\Biggr(-(2k+1)\pi+\operatorname{Arg}{\biggr(1+\frac{i(2k+1)}{8y_{3}}h^{\beta}\biggr)}-\operatorname{Im}{R_{k}}\Biggr)\\ \operatorname{Im}{z_{k}}=\frac{h}{2y_{3}}\Biggr(-\ln{\Bigl(8\pi y_{3}h^{-\beta}\Bigr)}-\ln{\left\lvert 1+\frac{i(2k+1)}{8y_{3}}h^{\beta}\right\rvert}+\operatorname{Re}{R_{k}}\Biggr)\end{cases}.$ (5.16)
@@
with $t=\frac{(2k+1)^{2}}{64y_{3}^{2}}h^{2\beta}$. Considering the imaginary part in Eq.˜5.16, we get
$\left\lvert\operatorname{Im}{z_{k}}+\frac{h}{2y_{3}}\ln{\Bigl(8\pi y_{3}h^{-\beta}\Bigr)}\right\rvert=\frac{h}{2y_{3}}\left\lvert\operatorname{Re}{R_{k}}-\ln{\left\lvert 1+\frac{i(2k+1)}{8y_{3}}h^{\beta}\right\rvert}\right\rvert$. (5.17)
@@
[…]